OpenStudy (canimcan):

Good morning friends! :) I need help with this equation [( ((xo-1)*(y-1))/(yo-1) + 1)^2 + 2y^2 +3((zo-1)(y-1))/(yo-1) + 1)]^2 =1. Thank you for sugestions!

OpenStudy (canimcan):

I need to fint decriminant

OpenStudy (will.h):

You may find the discriminant using B^2 -4ac

OpenStudy (canimcan):

Result is D=4[2((xo-1)/(yo-1))^2 -3((xo-1)*(zo-1))/(yo-1)*(yo-1) +3((zo-1)/(yo-1))^2 -2(xo-1)/(yo-1) -6(zo-1)/(yo-1)+3]=0 How to get it. Yes i am using b^2-4ac

OpenStudy (zzr0ck3r):

This is super confusing, the discriminant is something in a quadratic formula, this is not quadratic. Is there another definition?

OpenStudy (kevin):

This is question about matrix?

OpenStudy (kevin):

It will be more easy if you upload the picture :D

OpenStudy (mathmate):

\([( ((xo-1)*(y-1))/(yo-1) + 1)^2 + 2y^2 +3((zo-1)(y-1))/(yo-1) + 1)]^2 =1\) There are mismatched parentheses in the above expression. Is it: \(\Large ((\frac{(xo-1)*(y-1)}{yo-1}+1)^2+2y^2 +\frac{3(zo-1)*(y-1)}{yo-1}+1)^2 =1\) if not, please edit/correct. Please, in the future, double-check complicated expressions before/after posting. Helper's who spend time decoding your expression cannot help other users. Also, if possible, write in LaTeX format so you can do a good verification.

OpenStudy (canimcan):

#Matmate It is 3*[(zo-1)*(y-1)/(yo-1) +1]^2 ,my mistake.

OpenStudy (canimcan):

Rest of expression is correct.

OpenStudy (mathmate):

Like this? \(\Large (\frac{(xo-1)*(y-1)}{yo-1}+1)^2+2y^2 +3(\frac{(zo-1)*(y-1)}{yo-1}+1)^2 =1\) It makes a little more sense, because now it is a quadratic in y.

OpenStudy (mathmate):

To help you expand the expression a little easier, substitute x1=x0-1, y1=y0-1,z1=z0-1.

OpenStudy (canimcan):

Yes it is like you wrote.Alright I'll substitute that.